# Revisiting the relaxation of constraints in gauge theories

### v2

@@ 1 Introduction
We find it interesting to add yet another aspect to this theme. The paper [5] is written in the Hamiltonian language or, one can say, in terms of the first-order action functionals. We show that, starting from the proper Lagrangian description of gauge theories, the relaxation happens [^1] As we will see below, it is only for those of them in which the ”gauge hits twice” [6]. when we take the primary constraints only, and fix the gauge by adding conditions which are of second class with them.
@@ 2 The simple case of electrodynamics in vacuum
Let us first discuss the most classical field theory, that of electrodynamics. For the sake of simplicity, we take it in vacuum. The dynamical variable is a vector field $A^{\mu}$ and the field strength is defined as
@@
Note though that the initial idea of Dirac was different. In the first paper [8] on this topic, instead of $=$ and $\approx$ signs, he used $\equiv$ and $=$ respectively. He considered arbitrary independent variations of $q$, $\dot{q}$, $p$ variables (coordinate, velocity and momentum) and named an equation weak if it was valid to the zeroth order only, i.e. if it was immediately violated by a generic variation, and strong if it was correct up to the first order, at least, or maybe even exact. For example, the Lagrangian function $L\equiv L(q,\dot{q})$ was assumed to be given as a strong equality, while the definition of momenta $p=\frac{\partial L}{\partial\dot{q}}$ as a weak one. Then, eight years later, in a paper [^3] It was published together with his Hamiltonian form of general relativity, see below. [9] of the same title but in another journal, he changed the notation to the familiar $=$ and $\approx$ one and claimed to have put ”the method in a more direct and practical form”. The new motivation for defining a weak equality was that it is just a reminder of the fact that any Poisson bracket must be calculated before using equations which restrict the values of coordinates and momenta. In reality, there is nothing new in that. For example, a derivative of a function must also be taken before evaluating it at a particular point.
@@
while both Lagrange multipliers are arbitrary. We see that the definition of momentum is lost in its spatial longitudinal part, and the quantity $\partial_{i}(F_{0i}-\pi_{i})$ becomes pure gauge. The usual equations outside the transverse sector, including the Gauß law, are not there unless we assume that the electric field is now given by $\pi_{i}$ while $F_{0i}$ has partly lost its physical meaning.
@@ 5 The case of gravity
Roughly speaking, we have got an effective fluid without any pressure or shear stress. It actually reminds us of mimetic gravity [15, 16]. It would be interesting to see how far this analogy could go. We are not sure that one should be happy [4] about such possible addition. In mimetic gravity, it can be seen as a problem. Indeed, a pressureless ideal fluid does generally develop caustic singularities. For a physical fluid, we just realise that the no-pressure, or an ideal-fluid approximation has failed. In mimetic gravity it becomes a fundamental feature though.
@@ 6 Discussion and Conclusions
If one immediately fixes a gauge condition of the relaxation type, i.e. puts a new constraint of second class with the primary constraint, then the system no longer needs any secondary constraint, for the time preservation of the given constraints is then ensured by specifying the values of Lagrange multipliers. In Lagrangian language, it is because we have removed the equation which used to be a constraint for the time evolution.
All in all, what happens in the relaxation procedure is that we choose to impose a gauge fixing condition before looking for the secondary constraint. Moreover, we insist on using an incomplete gauge fixing thus producing a remnant gauge freedom, and also some fully new dynamics if the condition was imposed before deriving the equations of motion. Being of the lower time-derivative order, the momentum-constrained variable has naturally got a time derivative of the gauge parameter in its gauge transformation, and therefore fixing the gauge in terms of such variable is always incomplete, whatever boundary condition we might have imposed at the spatial infinity.
Gauge symmetry is an important concept in theoretical physics. Among other things, it allows us to write a nice action principle for equations of interaction fields. Therefore, removing it at the foundational level is a strange idea, to start with, let alone doing it with a gauge fixing which is necessarily incomplete even in perturbation theory. At the same time, an absolutely full gauge fixing would be difficult to naturally define in a field theory, and globally it is just impossible [21] for physically interesting non-Abelian theories due to Gribov ambiguity [^6] This is a topological obstruction [21]. Gribov himself found it [22] as a special case of remaining gauge freedom after imposing the Lorenz condition $\partial_{\mu}A_{a}^{\mu}=0$ in Euclidean time with boundary conditions at infinity..
In conclusion, we stress it again that the constraint relaxation idea [1, 2, 3, 4] is a way of modifying the classical theory, and different versions of it can be found in research papers over many years [5]. However, we find it curious to see that it is precisely what happens when we decide to fix a gauge before finding secondary constraints and do it in a second-class way with the primary constraints. Therefore, it is similar to fixing a gauge in the approach with an extended Hamiltonian, though done at another step of the Hamiltonian analysis.

### v3

@@ 1 Introduction
We find it interesting to add yet another aspect to this theme. The paper [5] is written in the Hamiltonian language or, one can say, in terms of the first-order action functionals. We show that, starting from the proper Lagrangian description of gauge theories, the relaxation happens [^1] As we will see below, it is only for those of them in which the ”gauge hits twice” [6]. when we take the primary constraints only and fix the gauge by adding conditions which are of second class with them.
@@ 2 The simple case of electrodynamics in vacuum
Let us first discuss the most classical field theory, that of electrodynamics. Its quantisation is very well-known, though the details can always be discussed [7]. For the sake of simplicity, we take it in vacuum. The dynamical variable is a vector field $A^{\mu}$ and the field strength is defined as
@@
Note though that the initial idea of Dirac was different. In the first paper [9] on this topic, instead of == and $\approx$ signs, he used $\equiv$ and == respectively. He considered arbitrary independent variations of $q$, $\dot{q}$, $p$ variables (coordinate, velocity and momentum) and named an equation weak if it was valid to the zeroth order only, i.e. if it was immediately violated by a generic variation, and strong if it was correct up to the first order, at least, or maybe even exact. For example, the Lagrangian function $L\equiv L(q,\dot{q})$ was assumed to be given as a strong equality, while the definition of momenta $p=\frac{\partial L}{\partial\dot{q}}$ as a weak one. Then, eight years later, in a paper [^3] It was published together with his Hamiltonian form of general relativity, see below. [10] of the same title but in another journal, he changed the notation to the familiar == and $\approx$ one and claimed to have put ”the method in a more direct and practical form”. The new motivation for defining a weak equality was that it is just a reminder of the fact that any Poisson bracket must be calculated before using equations which restrict the values of coordinates and momenta. In reality, there is nothing new in that. For example, a derivative of a function must also be taken before evaluating it at a particular point.
@@
while both Lagrange multipliers are arbitrary. We see that the definition of momentum is lost in its spatial longitudinal part, and the quantity $\partial_{i}(F_{0i}-\pi_{i})$ becomes pure gauge. The usual equations outside the transverse sector, that is the Gauß law, are not there unless we assume that the electric field is now given by $\pi_{i}$ while $F_{0i}$ has partly lost its physical meaning.
@@ 5 The case of gravity
Roughly speaking, we have got an effective fluid without any pressure or shear stress. It actually reminds us of mimetic gravity [18, 19]. It would be interesting to see how far this analogy could go. We are not sure that one should be happy [4] about such possible addition. In mimetic gravity, it can be seen as a problem. Indeed, a pressureless ideal fluid does generally develop caustic singularities. For a physical fluid, we just realise that the no-pressure, or an ideal-fluid approximation has failed. In mimetic gravity it becomes a fundamental feature though. Note in passing that we do not consider numerous modifications of mimetic gravity [20].
@@ 6 Discussion and Conclusions
If one immediately fixes a gauge condition of the relaxation type, i.e. puts a new constraint of second class with the primary constraint, then the system no longer needs any secondary constraint, for the time preservation of the given constraints is then ensured by specifying the values of Lagrange multipliers. In Lagrangian language, it is because we have removed the equation which used to be a constraint for the time evolution. Recently, the very same proposal, in the context of quantum cosmology, has appeared in a different but equivalent formulation [23].
All in all, what happens in the relaxation procedure is that we choose to impose a gauge fixing condition before looking for the secondary constraint. Moreover, we insist on using an incomplete gauge fixing thus producing a remnant gauge freedom, and also some fully new dynamics if the condition was imposed before deriving the equations of motion. Being of the lower time-derivative order, the momentum-constrained variable has naturally got a time derivative of the gauge parameter in its gauge transformation, and therefore fixing the gauge in terms of such variable is always incomplete, whatever boundary condition we might have imposed at the spatial infinity. Hence, it is not really correct to treat the lapse $N$, or the vector potential $A_{0}$ component, as merely a gauge parameter [23, 24].
Gauge symmetry is an important concept in theoretical physics. Among other things, it allows us to write a nice action principle for equations of interaction fields. Therefore, removing it at the foundational level is a strange idea, to start with, let alone doing it with a gauge fixing which is necessarily incomplete even in perturbation theory. At the same time, an absolutely full gauge fixing would be difficult to naturally define in a field theory, and globally it is just impossible [25] for the physically interesting non-Abelian theories due to Gribov ambiguity [^6] This is a topological obstruction [25]. Gribov himself found it [26] as a special case of remaining gauge freedom after imposing the Lorenz condition $\partial_{\mu}A_{a}^{\mu}=0$ in Euclidean time with boundary conditions at infinity..
In conclusion, we stress it again that the constraint relaxation idea [1, 2, 3, 4] is a way of modifying the classical theory, and different versions of it can be found in research papers over many years [5]. However, we find it curious to see that it is precisely what happens when we decide to fix a gauge before finding secondary constraints and do it in a second-class way with the primary constraints. Therefore, it is similar to fixing a gauge in the approach with an extended Hamiltonian, done at another step of the Hamiltonian analysis though.